# Prospect of Precision Cosmology and Testing General Relativity using Binary Black Holes- Galaxies Cross-correlation

### v1

@@ 4.1 Mathematical Formulation of Cross-Correlation Technique
The redshift information of GW binary source is derived by leveraging the cross-correlation of galaxy surveys with dark sirens. The expression for the relevant probability distribution is given by
@@
where $\rm{\vec{\vartheta}_{\text{GW}}=(D_{l}^{\text{GW}},\theta_{\text{GW}},\phi_{\text{GW}})}$ represents the GW data vector and $\rm{\vec{d}_{g}=(\delta_{g}(z_{g},\theta_{g},\phi_{g}))}$ denotes the galaxy data vector. Here, $\rm{\mathcal{P}(\vec{d}_{g}|P{gg}^{ss}(\vec{k},z))}$ is the posterior distribution of the galaxy density field and $\rm{\mathcal{L}(\vec{\vartheta}_{\text{GW}}|P^{ss}_{gg}(\vec{k},z),\vec{d}_{g}(z))}$ is the likelihood function. The superscript ’s’ stands for a spectroscopic (or photometric) survey that observes galaxies in this redshift space, which includes both the cosmological redshift and the additional redshift caused by the peculiar velocities of galaxies. The posterior distribution of the galaxy density field, conditional on the galaxy power spectrum, is described by the equation
@@
where $\rm{P^{ss}_{gg}(\vec{k},z)=b^{2}_{g}(k,z)(1+\beta_{g}\mu^{2}{\hat{k}})^{2}P_{m}(k,z)}$ is the three-dimensional power spectrum of galaxies, $\rm{b_{g}(k,z)}$ is the galaxy bias parameter, $\rm{\beta_{g}=\frac{f}{b_{g}}}$ with $\rm{f\equiv\frac{d\ln D}{d\ln a}}$ describing the logarithmic growth function, $\rm{\mu_{k}=\cos(\hat{n}\cdot\hat{k})}$ is the cosine of the angle between the line of sight $\rm{\hat{n}}$ and the Fourier modes $\rm{\hat{k}}$, and $\rm{n_{g}(z)=\frac{N_{g}(z)}{V_{s}}}$ is the number density of galaxies within the redshift bin $\rm{z}$. The Fourier transform of the galaxy distribution, $\rm{\delta^{s}_{g}(\vec{k},z)}$, is computed as $\rm{\int d^{3}\vec{r},\delta_{g}(\vec{r})e^{i\vec{k}\cdot\vec{r}}}$. Following the description of the probability distribution and posterior distribution, the likelihood function, $\rm{\mathcal{L}(\vec{\vartheta}_{\text{GW}}|P{gg}(\vec{k},z),\Theta_{n},\vec{d}_{g}(z))}$, is defined in Equation 25
[…]
where $\rm{\hat{P}(\vec{k},z)=\delta_{g}(\vec{k},z)\delta_{\text{GW}}^{*}(\vec{k},\Delta\Omega_{\text{GW}})}$ represents the cross power spectrum between the galaxy and GW data. The term $\rm{n_{\text{GW}}(z)=\frac{N_{\text{GW}}(D^{i}_{l}(z))}{V_{s}}}$ is the number density of gravitational wave sources described in terms of the number of objects within a specific luminosity distance bin. $\rm{V_{s}}$ denotes the total sky volume, and $\rm{P^{rr}_{\text{GW,GW}}(\vec{k},z)=b^{2}_{\text{GW}}(k,z)P_{m}(k,z)}$ with $\rm{b_{\text{GW}}(k,z)}$ being the GW bias parameter. This formulation for the likelihood function integrates cross power spectrum analysis, crucial for determining the correlation between galaxy distribution and GW sources, with a model that considers the biases and noise properties of both galaxies and gravitational waves. The exponential function captures the discrepancy between the observed and theoretical power spectra, reflecting the statistical nature of the inference process in cosmological studies.
@@ 4.2 Redshift Inference Using Cross-Correlation
Each of these surveys has its own advantages and disadvantages. DESI provides more accurate redshift measurements but lacks high-redshift coverage, whereas LSST, with its capability to observe high-redshift galaxies, complements DESI’s limitations. By cross-correlating these galaxy surveys with GW sources, we infer redshift information for dark sirens. DESI provides accurate measurements at low redshifts, while LSST allows us to extend this technique to higher redshifts. In this study, we explore both current and future GW detectors. Current detectors are limited in their ability to explore high redshifts, making DESI more effective for accurate measurements at these lower redshifts. However, future detectors will be capable of measuring sources at significant redshifts. LSST’s wide area coverage and deep field capabilities provide comprehensive measurements at high redshifts, while the precise redshifts from DESI enhance the accuracy of our low-redshift measurements. This study presents results from both surveys, demonstrating their applicability to both current and future GW detectors.
@@ 5 Baryon Acoustic Oscillation Scale from galaxy power spectrum
Several observatories are set to measure the BAO scale with remarkable precision, exploring the universe’s redshift depths. Among these instruments, DESI (Adame et al., 2023) is notable as a pioneering 5-year ground-based experiment designed to study BAO and the evolution of cosmic structures through redshift-space distortions. Spanning a substantial area of 14,000 square degrees in the sky, DESI aims to achieve highly accurate measurements of the BAO feature. The objective is to attain a precision of 0.5%, targeting the redshift range $\rm{0.0<z<3.7}$ (Adame et al., 2023). Additionally, Euclid (Laureijs et al., 2011) and the Vera Rubin Observatory (Abell et al., 2009; Ivezić et al., 2019), each covering an extensive sky area of 18,000 square degrees, will also be capable of making precise BAO measurements up to high redshifts. Figure 6 illustrates how the quantity $\rm{\theta^{2}w(\theta)}$ varies with angular separation at a specific redshift $\rm{z=0.275}$. The notable peak at $\theta\sim 7.0$ is a key observation, highlighting the characteristic scale associated with BAO, which is essential for our study. These calculations are based on the non-linear matter power spectrum for both DESI and LSST surveys, incorporating the selection functions specific to each survey as defined in Equation 29. As expected, the BAO feature is more pronounced for DESI than for LSST, reflecting the higher precision of the spectroscopic survey compared to the photometric one. This plot effectively illustrates the differences in observational capabilities and their impact on BAO feature detection between the two surveys. The covariance of $\rm{w(\theta)}$, denoted as $\rm{Cov_{\theta\theta^{\prime}}}$, is modeled accordingly.
@@
Here, $\rm{1/n}$ represents the shot noise, which is associated with the number density of galaxies $n$ or, more specifically, the number of objects per steradian. Additionally, $\rm{f_{sky}}$ indicates the fraction of the sky covered by the survey or observation (Cabré et al., 2007; Dodelson & Schmidt, 2020). Figure 7 shows how the covariance matrix, derived from the non-linear matter power spectrum, varies with angular separations at a specific redshift. The increase in the covariance matrix as both $\rm{\theta}$ and $\rm{\theta^{\prime}}$ increase suggests a correlation between these angular scales. Figure 8 offers a detailed view of how the BAO scale and its associated error (calculated using Equation 26) evolve with redshift for both DESI and LSST, highlighting the impact of redshift inaccuracies. This also illustrates that the uncertainty in BAO measurements is greater for LSST than for DESI, as anticipated.
@@ 6.1 Case with the fixed Hubble constant
This equation serves as the cornerstone for the analysis of the frictional component in cosmological studies, correlating three distinct length scales at a redshift ($\rm{z}$):
The measurement of redshift utilizes the cross-correlation technique between GW sources and galaxy surveys. The GW luminosity distance ($\rm{D_{l}^{GW}}$) for detectable GW events is determined using Bilby. The BAO scale $\rm{(\theta_{\text{BAO}})}$ is derived from the 2PACF using Equation 26. This method is versatile, allowing for application across various surveys without being restricted to any specific one. The sound horizon is measured through the analysis of acoustic oscillations in the CMB radiation, employing several missions (such as WMAP (Hinshaw et al., 2013), Planck (Aghanim et al., 2020), ACTPol (Thornton et al., 2016), and SPT-3G (Sobrin et al., 2022)). The first peak in the angular power spectrum of the CMB corresponds to the sound horizon scale at the time of recombination. The sound horizon value at the drag epoch, measured with high precision by Planck, is $\rm{147.09\pm 0.26}$ Mpc (Aghanim et al., 2020).
@@
[…]
$\times$ $dz^{i}\,P(r_{s})P(\theta_{\mathrm{BAO}}(z))P(z^{i})\mathcal{L}(D_{l}^{\mathrm{GW}^{i}}|\mathcal{F}(z^{i}),\theta_{\mathrm{BAO}}(z),r_{s},z^{i}).$
here, $\rm{P(\mathcal{F}(z)|\{D_{l}^{\mathrm{GW}}\},\{z\})}$ denotes the posterior probability density function of $\rm{\mathcal{F}(z)}$, given the data from GW sources $\{D_{l}^{\mathrm{GW}}\}$ and their redshifts $\rm{z}$, obtained using cross-correlation techniques between GW sources and galaxy surveys. The measurements of the BAO scale, $\rm{\theta_{\mathrm{BAO}}}$, at the redshift $\rm{z}$ of the GW source, are acquired from the galaxy surveys. The likelihood function is represented by $\rm{\mathcal{L}(D_{l}^{\mathrm{GW}^{i}}|\mathcal{F}(z^{i}),\theta_{\mathrm{BAO}}(z),r_{s},z^{i})}$. The terms $\rm{P(r_{s}),P(z^{i})}$, and $\rm{P(\theta_{\mathrm{BAO}}(z))}$ denote the prior probability distributions for the sound horizon $\rm{r_{s}}$, redshift $\rm{z}$, and BAO scale, respectively. The term $\rm{\Pi(\mathcal{F}(z))}$ represents the prior distribution for $\rm{\mathcal{F}(z)}$, which reflects our prior knowledge or assumptions about the frictional term before incorporating the measured data. A flat prior is adopted on $\rm{\mathcal{F}(z)}$, ranging from 0.1 to 2.
@@ 6.3 Case with varying the Hubble constant
In the preceding sections, our analysis primarily focused on evaluating the frictional term denoted as $\rm{\mathcal{F}(z)}$ in Equation 31, assuming a fixed Hubble constant ($\rm{H_{0}}$). However, in this section, we jointly infer both $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$. This extension enables the simultaneous estimation of both the frictional term $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$, derived from sound horizon measurements that remain constant across redshift, serving as an overall normalization on $\rm{\mathcal{F}(z)}$ measurements.
Employing a hierarchical Bayesian analysis, our unified framework efficiently estimates these two parameters by incorporating diverse observational datasets, including GW measurements, CMB measurements, and large-scale structure surveys, as discussed in previous sections. The hierarchical organization of these datasets ensures robust integration, effectively handling their unique uncertainties and systematics. The resulting correlated constraints significantly contribute to our understanding of cosmological phenomena, addressing potential discrepancies between datasets and revealing insightful connections among fundamental cosmological parameters. The equation representing the joint posterior distribution on $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$ for these sources is given as in Equation 34.
[…]
In this equation, $\rm{P(\mathcal{F}(z),H_{0}|\{D_{l}^{GW}\},\{z\})}$ denotes the joint posterior probability density function of $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$. The term $\rm{\mathcal{L}(D_{l}^{GW^{i}}|\mathcal{F}(z)^{i},H_{0}^{i},\theta_{BAO}(z),r_{s},z^{i})}$ corresponds to the likelihood function. The term $\rm{\Pi(H_{0})}$ denotes the prior distribution for $\rm{H_{0}}$, encapsulating our knowledge or assumptions regarding the Hubble constant before incorporating measured data. All other terms represent the same entities as defined in Section 6.1. In our hierarchical Bayesian framework, we adopt a flat prior assumption, indicating a lack of prior knowledge about the parameters for both the frictional term ($\mathcal{F}(z)$) and the cosmic expansion rate ($\rm{H_{0}}$). The parameter ranges are defined from 0.1 to 2.0 on $\rm{\mathcal{F}(z)}$ and from 20 km/s/Mpc to 120 km/s/Mpc on $\rm{H_{0}}$.

### v2

@@ 4.1 Mathematical Formulation of Cross-Correlation Technique
To estimate the redshift ($\rm{z}$) of a GW binary source when its EM is not detectable, we utilize the cross-correlation between galaxy surveys and dark sirens. The relevant probability distribution for this estimation is given by:
@@
In this expression, $\rm{\vec{\vartheta}_{\text{GW}}=(D_{l}^{\text{GW}},\theta_{\text{GW}},\phi_{\text{GW}})}$ represents the GW data vector, which includes the luminosity distance $\rm{D_{l}^{GW}}$ to the GW source, as well as the sky localization of the source $(\theta_{\text{GW}},\phi_{\text{GW}})$. The galaxy data vector $\rm{\vec{d}_{g}=(\delta_{g}(z_{g},\theta_{g},\phi_{g}))}$ includes the redshift information of the galaxy $\rm{z_{g}}$ as well as the sky position $(\rm{\theta_{g},\phi_{g}})$. The term $\rm{\mathcal{L}(\vec{\vartheta}_{\text{GW}}|P^{ss}_{gg}(\vec{k},z),\vec{d}_{g}(z))}$ represents the likelihood function, which measures how well the GW data fits the galaxy density field based on a model for the galaxy power spectrum $\rm{P^{ss}_{gg}(\vec{k},z)}$. This power spectrum describes the distribution of galaxy densities in redshift space, accounting for both cosmological redshift and additional redshift effects due to the peculiar velocities of galaxies. The term $\rm{\mathcal{P}(\vec{d}_{g}|P{gg}^{ss}(\vec{k},z))}$ denotes the posterior distribution of the galaxy density field given the power spectrum $\rm{P^{ss}_{gg}(\vec{k},z)}$. It provides the probability of observing the galaxy data under the assumed model for the galaxy power spectrum. The superscript ’s’ indicates that the galaxy survey in question is either a spectroscopic or photometric survey, observing galaxies in redshift space, which includes both cosmological redshift and redshift due to peculiar velocities. Overall, the probability distribution, $\rm{\mathcal{P}(z|\vec{\vartheta}_{GW},\vec{d}_{g})}$ combines information from GW data and galaxy surveys to estimate the redshift of the GW source. The posterior distribution of the galaxy density field, conditional on the galaxy power spectrum, is described by the equation
@@
where $\rm{P^{ss}_{gg}(\vec{k},z)=b^{2}_{g}(k,z)(1+\beta_{g}\mu^{2}{\hat{k}})^{2}P_{m}(k,z)}$ is the three-dimensional power spectrum of galaxies, $\rm{b_{g}(k,z)}$ is the galaxy bias parameter, $\rm{\beta_{g}=\frac{f}{b_{g}}}$ with $\rm{f\equiv\frac{d\ln D}{d\ln a}}$ describing the logarithmic growth function, $\rm{\mu_{k}=\cos(\hat{n}\cdot\hat{k})}$ is the cosine of the angle between the line of sight $\rm{\hat{n}}$ and the Fourier modes $\rm{\hat{k}}$, and $\rm{n_{g}(z)=\frac{N_{g}(z)}{V_{s}}}$ is the number density of galaxies within the redshift bin $\rm{z}$. The Fourier transform of the galaxy distribution, $\rm{\delta^{s}_{g}(\vec{k},z)}$, is computed as $\rm{\int d^{3}\vec{r},\delta_{g}(\vec{r})e^{i\vec{k}\cdot\vec{r}}}$. Following the description of the probability distribution and posterior distribution, the likelihood function, $\rm{\mathcal{L}(\vec{\vartheta}_{\text{GW}}|P{gg}(\vec{k},z),\Theta_{n},\vec{d}_{g}(z))}$, is defined in Equation(25)
$\mathrm{\mathcal{L}\propto\exp\bigg{(}-\frac{V_{s}}{4\pi^{2}}\int k^{2}dk\int d\mu_{k}\frac{\bigg{(}\hat{P}(\vec{k},\Delta\Omega_{GW})-b_{g}(k,z)b_{GW}(k,z)(1+\beta_{g}\mu_{\hat{k}}^{2})P_{m}(k,z)e^{-\frac{k^{2}}{k^{2}_{\rm eff}}}\bigg{)}^{2}}{2\big{(}P^{ss}_{gg}(\vec{k},z)+n_{g}(z)^{-1}\big{)}\big{(}P^{rr}_{GW\,GW}(\vec{k},z)+n_{GW}(z)^{-1}\big{)}}\bigg{)},}$ (25)
where $\rm{\hat{P}(\vec{k},z)=\delta_{g}(\vec{k},z)\delta_{\text{GW}}^{*}(\vec{k},\Delta\Omega_{\text{GW}})}$ represents the cross power spectrum between the galaxy and GW data. The term $\rm{n_{\text{GW}}(z)=\frac{N_{\text{GW}}(D^{i}_{l}(z))}{V_{s}}}$ is the number density of gravitational wave sources described in terms of the number of objects within a specific luminosity distance bin. $\rm{V_{s}}$ denotes the total sky volume, and $\rm{P^{rr}_{\text{GW,GW}}(\vec{k},z)=b^{2}_{\text{GW}}(k,z)P_{m}(k,z)}$ with $\rm{b_{\text{GW}}(k,z)}$ being the GW bias parameter. This formulation for the likelihood function integrates cross power spectrum analysis, crucial for determining the correlation between galaxy distribution and GW sources, with a model that considers the biases and noise properties of both galaxies and gravitational waves. The exponential function captures the discrepancy between the observed and theoretical power spectra, reflecting the statistical nature of the inference process in this studies due to the sky localization error.
@@ 4.2 Redshift Inference Using Cross-Correlation
Each galaxy survey has complementary advantages: DESI offers precise spectroscopic redshift measurements but is limited to lower redshifts, while LSST extends observations to higher redshifts with photometric measurements. By cross-correlating these surveys with GW sources, we can infer redshifts for dark sirens. DESI is particularly effective for low-redshift measurements with current GW detectors, whereas LSST’s wide coverage and high-redshift reach are better suited for future detectors. This study highlights the complementary roles of DESI and LSST in enhancing redshift inference across different GW detector capabilities.
@@ 5 Baryon Acoustic Oscillation Scale from galaxy power spectrum
Several observatories are set to measure the BAO scale with remarkable precision, exploring the universe’s redshift depths. Among these instruments, DESI (Adame et al., 2023) is notable as a pioneering 5-year ground-based experiment designed to study BAO and the evolution of cosmic structures through redshift-space distortions. Spanning a substantial area of 14,000 square degrees in the sky, DESI aims to achieve highly accurate measurements of the BAO feature. The objective is to attain a precision of 0.5%, targeting the redshift range $\rm{0.0<z<3.7}$ (Adame et al., 2023). Additionally, Euclid (Laureijs et al., 2011) and the Vera Rubin Observatory (Abell et al., 2009; Ivezić et al., 2019), each covering an extensive sky area of 18,000 square degrees, will also be capable of making precise BAO measurements up to high redshifts. Figure 6 illustrates how the quantity $\rm{\theta^{2}w(\theta)}$ varies with angular separation at a specific redshift $\rm{z=0.275}$. The notable peak at $\theta\sim 7.0$ is a key observation, highlighting the characteristic scale associated with BAO, which is essential for our study. These calculations are based on the non-linear matter power spectrum for both DESI and LSST surveys, incorporating the selection functions specific to each survey as defined in Equation (29). As expected, the BAO feature is more pronounced for DESI than for LSST, reflecting the higher precision of the spectroscopic survey compared to the photometric one. This plot effectively illustrates the differences in observational capabilities and their impact on BAO feature detection between the two surveys. The covariance of $\rm{w(\theta)}$, denoted as $\rm{Cov_{\theta\theta^{\prime}}}$, is modeled accordingly.
@@
Here, $\rm{1/n}$ represents the shot noise, which is associated with the number density of galaxies $n$ or, more specifically, the number of objects per steradian. Additionally, $\rm{f_{sky}}$ indicates the fraction of the sky covered by the survey or observation (Cabré et al., 2007; Dodelson & Schmidt, 2020). Figure 7 shows how the covariance matrix, derived from the non-linear matter power spectrum, varies with angular separations at a specific redshift. The increase in the covariance matrix as both $\rm{\theta}$ and $\rm{\theta^{\prime}}$ increase suggests a correlation between these angular scales. Figure 8 offers a detailed view of how the BAO scale and its associated error (calculated using Equation (26)) evolve with redshift for both DESI and LSST, highlighting the impact of redshift inaccuracies. This also illustrates that the uncertainty in BAO measurements is greater for LSST than for DESI, as anticipated.
@@ 6.1 Case with the fixed Hubble constant
This equation serves as the cornerstone for the analysis of the frictional component in cosmological studies, correlating three distinct length scales at a redshift ($\rm{z}$). The measurement of redshift utilizes the cross-correlation technique between GW sources and galaxy surveys. The GW luminosity distance ($\rm{D_{l}^{GW}}$) for detectable GW events is determined using Bilby. The BAO scale $\rm{(\theta_{\text{BAO}})}$ is derived from the 2PACF using Equation (26). This method is versatile, allowing for application across various surveys without being restricted to any specific one. The sound horizon is measured through the analysis of acoustic oscillations in the CMB radiation, employing several missions (such as WMAP (Hinshaw et al., 2013), Planck (Aghanim et al., 2020), ACTPol (Thornton et al., 2016), and SPT-3G (Sobrin et al., 2022)). The first peak in the angular power spectrum of the CMB corresponds to the sound horizon scale at the time of recombination. The sound horizon value at the drag epoch, measured with high precision by Planck, is $\rm{147.09\pm 0.26}$ Mpc (Aghanim et al., 2020).
@@
$\mathrm{P(\mathcal{F}(z))\propto\Pi(\mathcal{F}(z))\prod_{i=1}^{n_{\mathrm{GW}}}\iiiint dr_{s}\,dD_{l}^{\mathrm{GW}^{i}}\,d\theta_{\mathrm{BAO}}(z^{i})}$ (32)
$\times$ $\mathrm{dz^{i}\,P(r_{s})P(\theta_{\mathrm{BAO}}(z^{i}))P(z^{i})\mathcal{L}(D_{l}^{\mathrm{GW}^{i}}|\mathcal{F}(z^{i}),\theta_{\mathrm{BAO}}(z^{i}),r_{s},z^{i}),}$
here, $\rm{P(\mathcal{F}(z))}$ denotes the posterior probability density function of $\rm{\mathcal{F}(z)}$, given the data from GW sources $\rm{\{D_{l}^{\mathrm{GW}}\}}$ and their redshifts $\rm{z}$, obtained using cross-correlation techniques between GW sources and galaxy surveys. The measurements of the BAO scale, $\rm{\theta_{\mathrm{BAO}}}$, at the redshift $\rm{z}$ of the GW source, are acquired from the galaxy surveys. The likelihood function is represented by $\rm{\mathcal{L}(D_{l}^{\mathrm{GW}^{i}}|\mathcal{F}(z^{i}),\theta_{\mathrm{BAO}}(z^{i}),r_{s},z^{i})}$. The terms $\rm{P(r_{s}),P(z^{i})}$, and $\rm{P(\theta_{\mathrm{BAO}}(z^{i}))}$ denote the prior probability distributions for the sound horizon $\rm{r_{s}}$, redshift $\rm{z}$, and BAO scale, respectively. The term $\rm{\Pi(\mathcal{F}(z))}$ represents the prior distribution for $\rm{\mathcal{F}(z)}$, which reflects our prior knowledge or assumptions about the frictional term before incorporating the measured data. A flat prior on $\rm{\mathcal{F}(z)}$ ranging from 0.1 to 2 was chosen to ensure a non-informative prior that covers a broad range of possible deviations from GR while avoiding extreme values that are physically implausible or poorly constrained by the data. This range allows us to explore a wide parameter space without imposing strong biases on the results.
@@ 6.3 Case with varying the Hubble constant
In the preceding sections, our analysis focused on evaluating the frictional term $\rm{\mathcal{F}(z)}$ from Equation (31), assuming a fixed Hubble constant ($\rm{H_{0}}$). However, in this section, we extend our analysis to jointly infer both $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$. This approach treats $\rm{H_{0}}$ as a free parameter, serving as an overall normalization for the $\rm{\mathcal{F}(z)}$ measurements based on the sound horizon, which remains constant across the relevant redshifts.
Our framework employs a hierarchical Bayesian analysis to jointly estimate $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$, effectively handling their unique uncertainties and systematics. The resulting correlated constraints significantly contribute to our understanding of cosmological phenomena, addressing potential discrepancies between datasets and revealing insightful connections among fundamental cosmological parameters. The equation representing the joint posterior distribution on $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$ for these sources is given as in Equation (34).
$\mathrm{P(\mathcal{F}(z),H_{0})\propto\Pi(\mathcal{F}(z))\Pi(H_{0})\prod_{i=1}^{n_{\mathrm{GW}}}\iiiint dr_{s}\,dD_{l}^{\mathrm{GW}^{i}}\,d\theta_{\mathrm{BAO}}(z^{i})\,dz^{i}P(r_{s})P(\theta_{\mathrm{BAO}}(z^{i}))P(z^{i})\mathcal{L}(D_{l}^{\mathrm{GW}^{i}}|\mathcal{F}(z^{i}),H_{0},\theta_{\mathrm{BAO}}(z^{i}),r_{s},z^{i}).}$ (34)
In this equation, $\rm{P(\mathcal{F}(z),H_{0})}$ denotes the joint posterior probability density function of $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$. The term $\rm{\mathcal{L}(D_{l}^{GW^{i}}|\mathcal{F}(z)^{i},H_{0}^{i},\theta_{BAO}(z^{i}),r_{s},z^{i})}$ corresponds to the likelihood function. The term $\rm{\Pi(H_{0})}$ denotes the prior distribution for $\rm{H_{0}}$, encapsulating our knowledge or assumptions regarding the Hubble constant before incorporating measured data. All other terms represent the same entities as defined in Section 6.1. In our hierarchical Bayesian framework, we adopt a flat prior assumption, indicating a lack of prior knowledge about the parameters for both the frictional term ($\mathcal{F}(z)$) and the cosmic expansion rate ($\rm{H_{0}}$). The parameter ranges are defined from 0.1 to 2.0 on $\rm{\mathcal{F}(z)}$ and from 20 km/s/Mpc to 120 km/s/Mpc on $\rm{H_{0}}$. The choice of prior $\rm{\mathcal{F}(z)}$ is taken large enough to ensure that it can capture theoretical predictions of beyond GR models(Deffayet & Menou, 2007; Saltas et al., 2014; Nishizawa, 2018a; Belgacem et al., 2018a, b; Lombriser & Taylor, 2016; Lombriser & Lima, 2017).